# Evaluating Generalization and Representation Stability in Small LMs via Prompting, Fine-Tuning and Out-of-Distribution Prompts

### v1

@@ 1 Introduction
Few-shot prompting and supervised fine-tuning are two widely adopted strategies for adapting pretrained language models (LMs) to downstream tasks. Prompting adapts models by conditioning on in-context examples at inference time without updating model parameters (Brown et al., 2020), whereas fine-tuning involves directly optimizing the model on labeled data. While prompting is attractive for its flexibility and efficiency, its reliability in low-resource settings—particularly for small-scale language models like GPT-2 (Radford et al., 2019) and DistilGPT2 (Sanh et al., 2019)—remains uncertain.
@@ 2 Scaling and Shot Count: Accuracy and Representation Geometry
However, a t-SNE visualization  (van der Maaten & Hinton, 2008) of the final hidden state embeddings from each model (see Appendix 8.2, Figure 3) reveals an interesting asymmetry: although gpt2-medium achieved the highest task classification accuracy, its prompt representations did not form the most visually distinct clusters in the 2D projection. In contrast, gpt2 (base) exhibited clearer task-level separation under t-SNE, suggesting that the embeddings of larger models may become more abstract or compressed, making them less amenable to low-dimensional geometric partitioning.
@@ 4 Generalization to Out-of-Distribution (OOD) Prompts
We investigate how model behavior changes under variations in prompt formulation, using three semantically aligned but structurally distinct formats: standard, QA-style, and good/bad. While the underlying task—binary sentiment classification on IMDb—remains fixed, these prompt types differ in phrasing and label tokens (Reynolds & McDonell, 2021). Prompt construction details are provided in Appendix 8.1.
@@
Prompting results, shown in Figure 2(a), reveal that all models maintain consistent accuracy between the standard and QA-style formats. However, performance drops sharply for the good/bad format, especially in gpt2-medium, suggesting strong dependence on lexical alignment (Webson & Pavlick, 2022). Interestingly, distilgpt2 is more stable, possibly due to reduced reliance on pretraining-specific surface cues.

### v2

@@ 1 Introduction
Few-shot prompting and supervised fine-tuning are two widely adopted strategies for adapting pretrained language models (LMs) to downstream tasks. Prompting adapts models by conditioning on in-context examples at inference time without updating model parameters Brown et al. (2020), whereas fine-tuning involves directly optimizing the model on labeled data. While prompting is attractive for its flexibility and efficiency, its reliability in low-resource settings—particularly for small-scale language models like GPT-2 Radford et al. (2019) and DistilGPT2 Sanh et al. (2019)—remains uncertain.
@@ 2 Scaling and Shot Count: Accuracy and Representation Geometry
However, a t-SNE visualization  van der Maaten & Hinton (2008) of the final hidden state embeddings from each model (see Appendix 8.2, Figure 3) reveals an interesting asymmetry: although gpt2-medium achieved the highest task classification accuracy, its prompt representations did not form the most visually distinct clusters in the 2D projection. In contrast, gpt2 (base) exhibited clearer task-level separation under t-SNE, suggesting that the embeddings of larger models may become more abstract or compressed, making them less amenable to low-dimensional geometric partitioning.
@@ 4 Generalization to Out-of-Distribution (OOD) Prompts
We investigate how model behavior changes under variations in prompt formulation, using three semantically aligned but structurally distinct formats: standard, QA-style, and good/bad. While the underlying task—binary sentiment classification on IMDb—remains fixed, these prompt types differ in phrasing and label tokens Reynolds & McDonell (2021). Prompt construction details are provided in Appendix 8.1.
@@
Prompting results, shown in Figure 2(a), reveal that all models maintain consistent accuracy between the standard and QA-style formats. However, performance drops sharply for the good/bad format, especially in gpt2-medium, suggesting strong dependence on lexical alignment Webson & Pavlick (2022). Interestingly, distilgpt2 is more stable, possibly due to reduced reliance on pretraining-specific surface cues.